# A chaotic lattice field theory in two dimensions

### v2

@@ III.2 A many-body example: deterministic $\phi^{4}$ theory
III.3 Field theories that are first order in time
@@
IV Spatiotemporal stability of a periodic state
@@
IV.1 Primitive cell stability
@@
As an example, consider a periodic state $c$ over the one-dimensional primitive cell of period n{n}, Sec. I.2. For a periodic state $\mathsf{\Phi}_{c}$ of periodicity $\mathbb{A}={n}$, the orbit Jacobian matrix is
@@
where $\lambda$ is the stability exponent per unit-lattice-volume, with the exact steady state expression given below in Eq. (108).
@@
where $\lambda_{c}$ is the Birkhoff average Eq. (19) of the logarithms of orbit Jacobian matrix’s eigenvalues. This is a generalization of the temporal periodic orbit Floquet (or ‘Lyapunov’) stability exponent per unit time to any multi-periodic state, in any spatiotemporal dimension. (Continued in Sec. IX.1.)
IV.2 Bravais lattice stability
@@
Next, an elementary but essential observation. Consider a period-3 field configuration Eq. (18) obtained by a translation of another period-3 field configuration in its orbit. Or a period-6 field configuration obtained by a repeat of a period-3 field configuration. The orbit Jacobian operator Eq. (69) for all these field configurations is the same, of period 3. So, as announced in Eq. (2), and elaborated in Sec. IX, its spectrum is a property of its orbit, irrespective of whether it is computed over a prime periodic state primitive cell, or any lager primitive cell tiled by repeats of a prime periodic state.
@@
As we show in Sec. IX, for infinite lattices the appropriate notion of stability is the stability exponent Eq. (68) per unit-lattice-volume, averaged over the first Brillouin zone, which we evaluate by means of the Floquet-Bloch theorem.
V Bravais lattices
Periodic orbit theory of a time-evolving dynamical system on a one-dimensional temporal lattice is organized by grouping orbits of the same period together. [18, 19, 45, 20, 16] For systems characterized by several translational symmetries, one has to take care of multiple periodicities, or, in parlance of crystallography, organize the periodic orbit sums by corresponding Bravais lattices[53] introduced here in Sec. I.1.
@@ V Bravais lattices
Already in the case of chaotic lattice field theory over one-dimensional temporal integer lattice $\mathbb{Z}$ there is a sufficient amount of group-theoretical detail to merit the stand-alone companion paper I, [16] which treats in detail the time reversal invariance for $G=\textrm{D}_{\infty}$ dihedral space group of translations and reflections. Here we focus only on the two-dimensional square lattice translations Eq. (70), as a full symmetry treatment would distract the reader from the main trust of the paper, the construction of the spatiotemporal zeta function (Sec. X).
V.1 Bravais lattices of the square lattice
In crystallography there are 5 Bravais lattices over a two-dimensional space. The square lattice Eq. (5) is one of them. For brevity, whenever we refer here to a ‘Bravais lattice’, we mean one of the infinity of ‘full rank sublattices of the square lattice’ [126] that we now describe.
Figure 2: (Color online) The intersections of the light grey lines -lattice sites $z\in\mathbb{Z}^{2}$- form the integer square lattice Eq. (6). (a) Translations of the primitive cell $\mathbb{A}={[3\!\times\!2]_{1}}$ spanned by primitive vectors $\mathbf{a}_{1}=(3,0)$ and $\mathbf{a}_{2}=(1,2)$ define the Bravais lattice $\mathcal{L}_{\mathbb{A}}$. (b) The primitive vectors $\mathbf{a}_{1}=(2,-2)$ and $\mathbf{a}_{2}=(-1,4)$ form a primitive cell $\mathbb{A}^{\prime}$} equivalent to (a) by a unimodular transformation. The intersections (red points) of either set of dashed lines form the same Bravais lattice $\mathcal{L}_{\mathbb{A}}=\mathcal{L}_{\mathbb{A}^{\prime}}$. The volume Eq. (11) of either primitive cell is ${N}_{\mathcal{L}}=6$, the number of integer lattice sites within the cell, with the tips of primitive vectors and tiles’ outer boundaries belonging to the neighboring tiles. Continued in Fig. 3.
@@ V.1 Bravais lattices of the square lattice
‘Hermite normal form’ corresponds to the modular parameter $\tau$ values in the fundamental domain. If the corresponding torus is visualised as a glueing of a unit square into a tube, $\tau_{2}$ parameterizes how the tube is stretched, and $\tau_{1}$ parameterizes how it is twisted before its edges are stitched together.
@@
and to the set of lattice sites within the primitive parallelogram $\mathbb{A}$ as its primitive cell. Notation ${L}\!\times\!{T}]_{S}$} refers to primitive cell being a rectangle of spatial width L{L}, temporal height T{T}, with the primitive cell above it shifted by $S$, see for example the $3\!\times\!2]_{1}$} primitive cell shown in Fig. 3 (b). In terms of lattice site fields, a field configuration $\phi_{z_{1}z_{2}}$ Eq. (9), $z_{1}z_{2}\in\mathbb{Z}^{2}$, satisfies the S{S}-corkscrew boundary condition [71],
@@
Figure 3: Examples of ${L}\!\times\!{T}]_{S}$} field configurations Eq. (74) or ‘bricks’, together with their spatiotemporal Bravais lattice tilings, visualized as brick walls. (a) $2\!\times\!1]_{1}$}, primitive vectors $\mathbf{a}_{1}=(2,0)$, $\mathbf{a}_{2}=(1,1)$; (b) $3\!\times\!2]_{1}$} of Fig. 2 (a), primitive vectors $\mathbf{a}_{1}=(3,0)$, $\mathbf{a}_{2}=(1,2)$. Rectangles enclose the primitive cell and its Bravais lattice translations. Continued in Fig. 5.
VI Orbits over two-dimensional lattices
For field theories studied here (Sec. III), the translation group $T$ Eq. (70) is a symmetry, as their defining equations Eq. (31) retain their form (are ‘equivariant’) under lattice translations. For square lattice, these are 2-dimensional translations of form $g=\mathsf{r}_{1}^{m_{1}}\mathsf{r}_{2}^{m_{2}}$. (For symmetries other than translations, see remarks at the beginning of Sec. V.)
Typically a translation operation acting on periodic state $\mathsf{\Phi}_{p}$ generates an equivalent (up to lattice sites relabelling) but state-space distinct periodic state $g\mathsf{\Phi}_{p}$. The totality of all actions of the translation group on periodic states foliates the state space into a union Eq. (16) of translational orbits
@@ VI Orbits over two-dimensional lattices
Bravais lattices $\mathcal{L}_{\mathbb{A}}$ (Sec. I.2) are infinite, and their translational symmetries Eq. (70) are infinite groups, but the orbit of a Bravais periodic state is finite, generated by the translations of the infinite lattice curled up into a ${N}_{\mathbb{A}}$-sites periodic primitive cell $\mathbb{A}$.
VI.1 Prime orbits over two-dimensional primitive cells
@@
It is easy to check whether a one-dimensional periodic state $\mathsf{\Phi}_{p}$ over a primitive cell $\mathbb{A}$ is prime, by comparing it to its translations, as in the period-6 example of Sec. I.3. We use this test as an operational definition of a prime periodic state over a primitive cell $\mathbb{A}$ for a hypercubic lattice in two (or any) dimensions.
@@
This notion of a ‘prime’ suffices to formulate our main result, the spatiotemporal zeta function (Sec. X.4) for field theories in two spacetime dimensions. However, we have to emphasize the implementing this for a given theory requires determination of all of its prime orbits, and that is hard problem, in a sense all of ‘chaos theory’. Here we use spatiotemporal cat to test the theory, but relegate details to Appendix B, and in the companion paper III [17] we apply the theory to several nonlinear field theories.
VI.2 Repeats of a prime orbit over two-dimensional primitive cells
Figure 4: (Color online) (a) Bravais lattice $\mathbb{A}={[6\!\times\!4]_{2}}$, blue dots, is a sublattice of Bravais lattice $\mathbb{A}_{p}={[3\!\times\!2]_{1}}$, blue and red dots. Its primitive cell $\mathbb{A}$} (green parallelogram spanned by primitive vectors (6,0) and (2,4)) is tiled by repeats of the primitive cell $\mathbb{A}_{p}$ (gray parallelogram spanned by primitive vectors (3,0) and (1,2)). The primitive vectors of the 2 Bravais lattices are related by $\mathbb{A}=\mathbb{A}_{p}\mathbb{R}$ where $\mathbb{R}={[2\!\times\!2]_{0}}$. (b) Transform the primitive cell $\mathbb{A}_{p}$ to the unit square of a new square lattice, where each unit square supports a multiplet of 6 fields belonging to a prime $\mathcal{L}_{\mathbb{A}_{p}}$-periodic state. In this new square lattice, the prime periodic state is a steady state whose primitive cell is a $1\!\times\!1]_{0}$ unit square (gray square), while the repeat of the prime is a $\mathcal{L}_{\mathbb{R}}$-periodic state, whose primitive cell is $\mathbb{R}={[2\!\times\!2]_{0}}$ (green square).
The simplest example of a prime periodic state is a steady state $\phi_{z}={\phi}$, invariant under all of system’s symmetries. Its primitive cell $1\!\times\!1]_{0}$} is the unit hypercube Eq. (5) of period-1 along every hypercube direction.
A periodic state obtained by tiling any larger primitive cell by repeats of steady state $\phi$} is not a prime periodic state. There is one such for each Bravais sublattice constructed in Sec. V.1, with $r_{1}$ copies of lattice site field $\phi{}$ horizontally, $r_{2}$ copies vertically, and tilt $0\leq s<r_{1}$ Eq. (74),
@@ VI.2 Repeats of a prime orbit over two-dimensional primitive cells
Next, note that every orbit is ‘steady’ in the sense that each orbit Eq. (78) is a fixed point of $T$, as any translation $g\,\mathcal{M}_{p}=\mathcal{M}_{p}$ only permutes the set of periodic states within the orbit, but leaves the set invariant. In particular (see Sec. IV.2), the stability of an orbit is its intrinsic, translation invariant ‘steady’ property.
A way to visualize this is by multiplying the Bravais lattice $\mathcal{L}_{\mathbb{A}_{p}}$ by $\mathbb{A}_{p}^{-1}$, sending it into the unit integer lattice, as in Fig. 4 (b): in other words, every Bravais lattice is a hypercubic lattice, under an appropriate change of coordinates. In this new integer lattice, the primitive cell $\mathbb{A}_{p}$ is the unit square that supports a multiplet of ${N}_{\mathbb{A}}$ periodic states belonging to the $\mathcal{L}_{\mathbb{A}_{p}}$ orbit. Under lattice translations, this multiplet is an ${N}_{\mathbb{A}}$-dimensional steady state.
@@
Figure 5: (Color online) (a) Bravais lattice $\mathbb{A}={[3\!\times\!2]_{1}}$ of Fig. 2, red dots, is a sublattice of Bravais lattice $\mathbb{A}^{\prime}={[3\!\times\!1]_{2}}$, blue and red dots, even though the primitive cell $\mathbb{A}$} (green parallelogram spanned by primitive vectors (3,0) and (1,2)) does not appear to be tiled by a repeat of the primitive cell $\mathbb{A}^{\prime}$} (blue parallelogram spanned by primitive vectors (3,0) and (2,1)). (b) If we shift the top edge of primitive cell $\mathbb{A}$} by 3 lattice units, to ${[3\!\times\!2]_{4}}={[3\!\times\!2]_{1}}$ (green parallelogram spanned by primitive vectors (3,0) and (4,2)), the tiling is clear.
A priori is not obvious that $3\!\times\!1]_{2}$} primitive cell tiles the $3\!\times\!2]_{1}$} primitive cell, Fig. 5 (a). But if you stack $3\!\times\!1]_{2}$} primitive cell in the shifted temporal direction by 2 then the left edge of the tile is shifted by 4 in the spatial direction. With the spatial period being 3, shift by 4 in the spatial direction is same as shift by 1. So the boundary conditions of $3\!\times\!2]_{1}$} primitive cell are satisfied by the repeat of the $3\!\times\!1]_{2}$} primitive cell.
@@
Figure 6: (Color online) Examples of spatiotemporal mosaic tilings Eq. (37) of $6\!\times\!6]_{0}$} primitive cell by repeats of smaller prime periodic states. (a) $3\!\times\!1]_{0}$} temporally steady state. (b) $1\!\times\!3]_{0}$} spatially steady state. (c) $2\!\times\!1]_{1}$} relative-periodic prime orbit, spatial period-2, temporal period-2; compare with Fig. 3 (a). (d) $3\!\times\!1]_{1}$} relative-periodic prime orbit, spatial period-3, temporal period-3. (e) $3\!\times\!2]_{0}$} spatial period-3, temporal period-2. (f) $3\!\times\!2]_{1}$} of figures 3 (b) and 5. It is a relative-periodic prime orbit, of spatial period-3, temporal period-6. See also Fig. 12 and Appendix A.
In summary: to determine all periodic states, it suffices to enumerate all Bravais lattices, then compute their prime orbits on their finite-dimensional primitive cells. Their stabilities, however, will have to be evaluated on the infinite Bravais lattices, as we shall show in Sec. IX.
VII Reciprocal lattice

### v3

@@ IV.2 A many-body example: deterministic $\phi^{4}$ theory
IV.3 Field theories that are first order in time
@@
V Spatiotemporal stability of a periodic state
@@
V.1 Primitive cell stability
@@
As an example, consider a periodic state $c$ over the one-dimensional primitive cell of period n{n}, Sec. II.2. For a periodic state $\mathsf{\Phi}_{c}$ of periodicity $\mathbb{A}={n}$, the orbit Jacobian matrix is
@@
where $\lambda$ is the stability exponent per unit-lattice-volume, with the exact steady state expression given in Eq. (108).
@@
where $\lambda_{c}$ is the Birkhoff average Eq. (19) of the logarithms of orbit Jacobian matrix’s eigenvalues. This is a generalization of the temporal periodic orbit Floquet (or ‘Lyapunov’) stability exponent per unit time to any multi-periodic state, in any spatiotemporal dimension. This will be discussed further in Sec. X.1.
V.2 Bravais lattice stability
@@
The next will be an elementary but essential observation. Consider a period-3 field configuration Eq. (18) obtained by a translation of another period-3 field configuration in its orbit. Or a period-6 field configuration obtained by a repeat of a period-3 field configuration. The orbit Jacobian operator Eq. (69) for all these field configurations is the same, of period 3. So, as announced in Eq. (2), and elaborated in Sec. X, its spectrum is a property of its orbit, irrespective of whether it is computed over a prime periodic state primitive cell, or any larger primitive cell tiled by repeats of a prime periodic state.
@@
As we show in Sec. X, for infinite lattices the appropriate notion of stability is the stability exponent Eq. (68) per unit-lattice-volume, averaged over the first Brillouin zone, which we evaluate by means of the Floquet-Bloch theorem.
VI Bravais lattices
Periodic orbit theory of a time-evolving dynamical system on a one-dimensional temporal lattice is organized by grouping orbits of the same period together. [18, 19, 45, 20, 16] For systems characterized by several translational symmetries, one has to take care of multiple periodicities, or, in parlance of crystallography, organize the periodic orbit sums by corresponding Bravais lattices[54] introduced here in Sec. II.1.
@@ VI Bravais lattices
Already in the case of chaotic lattice field theory over one-dimensional temporal integer lattice $\mathbb{Z}$ there is a sufficient amount of group-theoretical detail to merit the stand-alone companion paper I, [16] which treats in detail the time-reversal invariance for a $G=\textrm{D}_{\infty}$ dihedral space group of translations and reflections. Here we focus only on the two-dimensional square lattice translations Eq. (70), as a full symmetry treatment would distract the reader from the main trust of the paper, the construction of the spatiotemporal zeta function (Sec. XI).
VI.1 Bravais lattices of the square lattice
In crystallography, there are five Bravais lattices over a two-dimensional space. The square lattice Eq. (5) is one of them. For brevity, whenever we refer here to a ‘Bravais lattice’, we mean one of the infinity of ‘full rank sublattices of the square lattice’ [132] that we now describe.
@@ VI.1 Bravais lattices of the square lattice
‘Hermite normal form’ corresponds to the modular parameter $\tau$ values in the fundamental domain. If the corresponding torus is visualized as glueing of a unit square into a tube, $\tau_{2}$ parameterizes how the tube is stretched, and $\tau_{1}$ parameterizes how it is twisted before its edges are stitched together.
@@
and to the set of lattice sites within the primitive parallelogram $\mathbb{A}$ as its primitive cell. Notation ${L}\!\times\!{T}]_{S}$} refers to primitive cell being a rectangle of spatial width L{L}, temporal height T{T}, with the primitive cell above it shifted by $S$, see, for example the $3\!\times\!2]_{1}$} primitive cell shown in Fig. 6 (b). In terms of lattice site fields, a field configuration $\phi_{z_{1}z_{2}}$ Eq. (9), $z_{1}z_{2}\in\mathbb{Z}^{2}$, satisfies the S{S}-corkscrew boundary condition [73],
@@
Figure 6: Examples of ${L}\!\times\!{T}]_{S}$} field configurations Eq. (74) or ‘bricks’, together with their spatiotemporal Bravais lattice tilings, visualized as brick walls. (a) $2\!\times\!1]_{1}$}, primitive vectors $\mathbf{a}_{1}=(2,0)$, $\mathbf{a}_{2}=(1,1)$; (b) $3\!\times\!2]_{1}$} of Fig. 2 (a), primitive vectors $\mathbf{a}_{1}=(3,0)$, $\mathbf{a}_{2}=(1,2)$. Rectangles enclose the primitive cell and its Bravais lattice translations. This will be discussed further in Fig. 8.
VII Orbits over two-dimensional lattices
For field theories studied here (Sec. IV), the translation group $T$ Eq. (70) is a symmetry, as their defining equations Eq. (31) retain their form (are ‘equivariant’) under lattice translations. For square lattice, these are 2-dimensional translations of the form $g=\mathsf{r}_{1}^{m_{1}}\mathsf{r}_{2}^{m_{2}}$. (For symmetries other than translations, see remarks at the beginning of Sec. VI.)
Typically a translation operation acting on periodic state $\mathsf{\Phi}_{p}$ generates an equivalent (up to lattice sites relabeling) but state-space distinct periodic state $g\mathsf{\Phi}_{p}$. The totality of all actions of the translation group on periodic states foliates the state space into a union Eq. (16) of translational orbits
@@ VII Orbits over two-dimensional lattices
Bravais lattices $\mathcal{L}_{\mathbb{A}}$ (Sec. II.2) are infinite, and their translational symmetries Eq. (70) are infinite groups, but the orbit of a Bravais periodic state is finite, generated by the translations of the infinite lattice curled up into a ${N}_{\mathbb{A}}$-site periodic primitive cell $\mathbb{A}$.
VII.1 Prime orbits over two-dimensional primitive cells
@@
It is easy to check whether a one-dimensional periodic state $\mathsf{\Phi}_{p}$ over a primitive cell $\mathbb{A}$ is prime, by comparing it to its translations, as in the period-6 example of Sec. II.3. We use this test as an operational definition of a prime periodic state over a primitive cell $\mathbb{A}$ for a hypercubic lattice in two (or any) dimensions.
@@
This notion of a ‘prime’ suffices to formulate our main result, the spatiotemporal zeta function (Sec. XI.4) for field theories in two spacetime dimensions. However, we have to emphasize that implementing this for a given theory requires determination of all of its prime orbits, and that is hard problem, in a sense all of ‘chaos theory’. Here we use spatiotemporal cat to test the theory, but relegate details to Appendix B, and in the companion paper III [17] we apply the theory to several nonlinear field theories.
VII.2 Repeats of a prime orbit over two-dimensional primitive cells
$\phi_{z}={\phi}$, invariant under all of system’s symmetries. Its primitive cell $1\!\times\!1]_{0}$} is the unit hypercube Eq. (5) of period-1 along every hypercube direction.
A periodic state obtained by tiling any larger primitive cell by repeats of steady state $\phi$} is not a prime periodic state. There is one such for each Bravais sublattice constructed in Sec. VI.1, with $r_{1}$ copies of lattice site field $\phi{}$ horizontally, $r_{2}$ copies vertically, and tilt $0\leq s<r_{1}$ Eq. (74),
@@ VII.2 Repeats of a prime orbit over two-dimensional primitive cells
Next, note that every orbit is ‘steady’ in the sense that each orbit Eq. (78) is a fixed point of $T$, as any translation $g\,\mathcal{M}_{p}=\mathcal{M}_{p}$ only permutes the set of periodic states within the orbit, but leaves the set invariant. In particular (see Sec. V.2), the stability of an orbit is its intrinsic, translation invariant ‘steady’ property.
Figure 7: (a) Bravais lattice $\mathbb{A}={[6\!\times\!4]_{2}}$, blue dots, is a sublattice of Bravais lattice $\mathbb{A}_{p}={[3\!\times\!2]_{1}}$, blue and red dots. Its primitive cell $\mathbb{A}$} (green parallelogram spanned by primitive vectors (6,0) and (2,4)) is tiled by repeats of the primitive cell $\mathbb{A}_{p}$ (gray parallelogram spanned by primitive vectors (3,0) and (1,2)). The primitive vectors of the two Bravais lattices are related by $\mathbb{A}=\mathbb{A}_{p}\mathbb{R}$ where $\mathbb{R}={[2\!\times\!2]_{0}}$. (b) Transform the primitive cell $\mathbb{A}_{p}$ to the unit square of a new square lattice, where each unit square supports a multiplet of six fields belonging to a prime $\mathcal{L}_{\mathbb{A}_{p}}$-periodic state. In this new square lattice, the prime periodic state is a steady state whose primitive cell is a $1\!\times\!1]_{0}$ unit square (gray square), while the repeat of the prime is a $\mathcal{L}_{\mathbb{R}}$-periodic state, whose primitive cell is $\mathbb{R}={[2\!\times\!2]_{0}}$ (green square).
The simplest example of a prime periodic state is a steady state A way to visualize this is by multiplying the Bravais lattice $\mathcal{L}_{\mathbb{A}_{p}}$ by $\mathbb{A}_{p}^{-1}$, sending it into the unit integer lattice, as in Fig. 7 (b): in other words, every Bravais lattice is a hypercubic lattice, under an appropriate change of coordinates. In this new integer lattice, the primitive cell $\mathbb{A}_{p}$ is the unit square that supports a multiplet of ${N}_{\mathbb{A}}$ periodic states belonging to the $\mathcal{L}_{\mathbb{A}_{p}}$ orbit. Under lattice translations, this multiplet is an ${N}_{\mathbb{A}}$-dimensional steady state.
@@
Figure 8: (a) Bravais lattice $\mathbb{A}={[3\!\times\!2]_{1}}$ of Fig. 2, red dots, is a sublattice of Bravais lattice $\mathbb{A}^{\prime}={[3\!\times\!1]_{2}}$, blue and red dots, even though the primitive cell $\mathbb{A}$} (green parallelogram spanned by primitive vectors (3,0) and (1,2)) does not appear to be tiled by a repeat of the primitive cell $\mathbb{A}^{\prime}$} (blue parallelogram spanned by primitive vectors (3,0) and (2,1)). (b) If we shift the top edge of primitive cell $\mathbb{A}$} by three lattice units, to ${[3\!\times\!2]_{4}}={[3\!\times\!2]_{1}}$ (green parallelogram spanned by primitive vectors (3,0) and (4,2)), the tiling is clear.
A priori is not obvious that the $3\!\times\!1]_{2}$} primitive cell tiles the $3\!\times\!2]_{1}$} primitive cell, Fig. 8 (a). However, if you stack the $3\!\times\!1]_{2}$} primitive cell in the shifted temporal direction by 2 then the left edge of the tile is shifted by 4 in the spatial direction. With the spatial period being 3, shift by 4 in the spatial direction is the same as shift by 1. Therefore, the boundary conditions of the $3\!\times\!2]_{1}$} primitive cell are satisfied by the repeat of the $3\!\times\!1]_{2}$} primitive cell.
@@
In summary, to determine all periodic states, it suffices to enumerate all Bravais lattices, then compute their prime orbits on their finite-dimensional primitive cells. Their stabilities, however, will have to be evaluated on the infinite Bravais lattices, as we shall show in Sec. X.
VIII Reciprocal lattice